# Efficient and high-performance routing of lattice-surgery paths on three-dimensional lattice

### v3

@@ A.3 Extension of Dijkstra projection to many-body cases
    Perform a two-body Dijkstra projection for logical qubits $Q_{1}$ and $Q_{2}$.
@@
    For each $i=3,4,\ldots,m$, perform a two-body Dijkstra projection between logical qubit $Q_{i}$ and the nearest $Z$-boundary of the existing even segment, and adjust the kink between them to be even.
@@ B.1 Quantum circuit representation
The first step is to synthesize a quantum circuit with a sequence of quantum gates defined with the form listed in Table 8.
@@ B.2 FTQC machine instructions
The unitary instructions UNIT_* perform logical Hadamard and phase gates acting on target cells [brown2017poking, fowler2018low]. These operations require a neighboring cell: the Hadamard gate requires any adjacent cell, while the phase gate requires one attached to the Z-boundary [beverland2022assessing]. A previously stored measurement value in C can be used to condition these operations.
@@ B.3 Translation from circuit forms to instructions
This section describes how to translate a quantum circuit representation into FTQC instructions without location symbols. Quantum gates H and S are directly replaced with UNIT_H and UNIT_S, respectively. We decompose CCX into (CX, T, UNIT_H), following known constructions [nielsen2010quantum]. Next, we decompose CX into (INIT_X, MEAS_ZZ, MEAS_XX, MEAS_Z) and conditional (X, Z) operations using the method in Ref. [fowler2018low]. This decomposition requires an ancillary logical qubit, so we assign one ancillary qubit symbol for each data logical qubit. We then decompose the Pauli rotations (RX, RY, RZ) into (UNIT_S, UNIT_H, T) using the Ross–Selinger decomposition [ross2014optimal]. The T gate is further decomposed via magic-state consumption using (MEAS_MAGIC, MEAS_X) and conditional (UNIT_S, X) gates, following Ref. [fowler2018low]. Finally, we remove Pauli unitary operations X, Y, and Z and replace them with XOR operations using the Pauli frame technique. After these steps, the quantum circuit is fully decomposed into INIT_*, MEAS_*, UNIT_*, and XOR instructions without location operands.
Next, we perform placement and routing [yoshioka2022hunting], which fills in the location operands (L) to make the executable compatible with a specific machine layout. For simplicity, we assume that all required logical qubits are allocated initially and never deallocated. In our benchmark, logical qubits are arranged on the pattern shown in Fig. 1. The allocation rule for each evaluation is detailed in the next section. Finally, we assign ancillary locations and paths for UNIT_* and MEAS_* instructions. While this process can be performed greedily, our method can be used to improve throughput.
@@ B.4 Instance synthesis and simplification used in this paper
Regarding the comparison in Sec. 4.3, we constructed an instruction sequence consisting of CX and MEAS_MAGIC for the compatibility with EDPC. Also, we adopted the settings of the EDPC algorithm in Ref. [beverland2022surface], which are outlined below. Logical qubits are assigned from top left to bottom right. It is assumed that the data cells have identical boundary orientations. Magic state factories are placed at the position $(i,j)$ such that $i$ and $j$ are even and satisfying any of the following: $i=0$, i=W1i=W-1, $j=0$, or j=W1j=W-1, where $W$ is the width of the qubit-plane cells. See Fig. 10 of Ref. [beverland2022surface] for illustrations. It is also assumed that a magic state factory produces a single magic state every code beat.

### v4

@@ A.3 Extension of Dijkstra projection to many-body cases
    Perform the two-body Dijkstra projection for logical qubits $Q_{1}$ and $Q_{2}$.
@@
    For each $i=3,4,\ldots,m$, perform the two-body Dijkstra projection between logical qubit $Q_{i}$ and the nearest $Z$-boundary of the existing even segment, and adjust the kink between them to be even.
@@ B.1 Quantum circuit representation
The first step is to synthesize a quantum circuit with a sequence of quantum gates defined in the form listed in Table 8.
@@ B.2 FTQC machine instructions
The unitary instructions UNIT_* perform logical Hadamard and phase gates acting on target cells [8, 12]. These operations require a neighboring cell: the Hadamard gate requires any adjacent cell, while the phase gate requires one attached to the Z-boundary [3]. A previously stored measurement value in C can be used to condition these operations.
@@ B.3 Translation from circuit forms to instructions
This section describes how to translate a quantum circuit representation into FTQC instructions without location symbols. Quantum gates H and S are directly replaced with UNIT_H and UNIT_S, respectively. We decompose CCX into (CX, T, UNIT_H), following known constructions [34]. Next, we decompose CX into (INIT_X, MEAS_ZZ, MEAS_XX, MEAS_Z) and conditional (X, Z) operations using the method in Ref. [12]. This decomposition requires an ancillary logical qubit, so we assign one ancillary qubit symbol for each data logical qubit. We then decompose the Pauli rotations (RX, RY, RZ) into (UNIT_S, UNIT_H, T) using the Ross–Selinger decomposition [37]. The T gate is further decomposed via magic-state consumption using (MEAS_MAGIC, MEAS_X) and conditional (UNIT_S, X) gates, following Ref. [12]. Finally, we remove Pauli unitary operations X, Y, and Z and replace them with XOR operations using the Pauli frame technique. After these steps, the quantum circuit is fully decomposed into INIT_*, MEAS_*, UNIT_*, and XOR instructions without location operands.
Next, we perform placement and routing [44], which fills in the location operands (L) to make the executable compatible with a specific machine layout. For simplicity, we assume that all required logical qubits are allocated initially and never deallocated. In our benchmark, logical qubits are arranged on the pattern shown in Fig. 1. The allocation rule for each evaluation is detailed in the next section. Finally, we assign ancillary locations and paths for UNIT_* and MEAS_* instructions. While this process can be performed greedily, our method can be used to improve throughput.
@@ B.4 Instance synthesis and simplification used in this paper
Regarding the comparison in Sec. 4.3, we constructed an instruction sequence consisting of CX and MEAS_MAGIC for compatibility with EDPC. Also, we adopted the settings of the EDPC algorithm in Ref. [4], which are outlined below. Logical qubits are assigned from top left to bottom right. It is assumed that the data cells have identical boundary orientations. Magic state factories are placed at the position $(i,j)$ such that $i$ and $j$ are even and satisfying any of the following: $i=0$, i=W1i=W-1, $j=0$, or j=W1j=W-1, where $W$ is the width of the qubit-plane cells. See Fig. 10 of Ref. [4] for illustrations. It is also assumed that a magic state factory produces a single magic state every code beat.